A car travels 100m as it accelerates from rest at 25m/s. Calculate the acceleration.


1
Expert's answer
2022-01-14T09:46:21-0500

"v_f^2=v_i^2+2aS"


Rearranging for a,


"a=\frac{v_f^2-v_i^2}{2S}=\frac{(25\frac{m}{s})^2-(0\frac{m}{s})^2}{2\times100\ m}=3.125\frac{m}{s^2}"


Answer: "3.125\frac{m}{s^2}"